Can you solve ln(x-4) + ln(x+1) = ln (x-8)
Answer by stanbon(48535)

solve ln(x-4) + ln(x+1) = ln (x-8)
---------------
ln[(x-4)(x+1)] = ln(x-8)
---
x^2-3x-4 = x-8
x^2-4x+4 = 0
(x-2)^2 = 0
x = 2
---
But x cannot be 2 for ln(x-4) would be the natural log of a negative,
which does not exist.
-----------------------
Conclusion: No solution
-----------------------
Cheers,
Stan H.